OpenStudy (anonymous):

What kind of property is used in 2x=2x ?

OpenStudy (lgbasallote):

looks like reflexive...although i always confuse reflexive and symmetric

OpenStudy (anonymous):

Omg. Its on these CST Review packet my teacher gave but I didnt even learn it. Only Distibutice, Assoc., Comm., and Identity. :/

OpenStudy (lgbasallote):

oh....well distributive is a(b+c) = ab + ac associative is (a+b) + c = a + (b+c) commutative is a+ b = b + a identity is a= a

OpenStudy (anonymous):

Ty.